OpenStudy (testing123guest):

Write a fraction that is equivalent to 2/3.

imqwerty (imqwerty):

just multiply the numerator and denominator both of 2/3 by some integer(except for 0 and 1) and you get your required fraction which is equivalent to 2/3

KjSaif (kjsaif):

2/3 hmmmmm 2/3 looking at this what do you think? (it's okay if you're wrong,were here to help!)

KjSaif (kjsaif):

what do you think is equivalent? with the given help from @imqwerty

OpenStudy (adelyn_helps):

Some equivalent fractions of 2/3 are: 2/3 = 4/6 = 6/9 = 8/12 = 10/15 = 12/18 = 14/21 = 16/24 = 18/27 = 20/30 = 22/33 = 24/36 = 26/39 = 28/42 = 30/45 = 32/48 = 34/51 = 36/54 = 38/57 = 40/60

OpenStudy (marypuro):

multiply the top and bottom both by the same number to get an equivalent fractions. So 2/3 *2/2= 4/6. just make sure once fraction is reduced it equals your 2/3